OpenStudy (anonymous):

PLEASE HELP! How do you solve What is the slope and y-intercept of the graph of 4x + 3y = 12? I tried solving it but I have no idea...

OpenStudy (anonymous):

3y = -4x + 12 y = -4/3x + 4 slope = -4/3 y intercept = 4

OpenStudy (anonymous):

thank you so much!